OpenStudy (anonymous):

Find the general term of the arithmetic sequence where a[4] = 95 a[7] = 110. ?

OpenStudy (anonymous):

a[4]=a+3d=95 and a[7]=a+6d=110 solve for both a and d

OpenStudy (anonymous):

i'm still confused.

OpenStudy (amistre64):

general term? is that the same as a common difference?

OpenStudy (amistre64):

or is that spose to be the first term of the seq?

OpenStudy (anonymous):

just solve the linear system to get the a and d. general term is a[n]=a+(n-1)d

OpenStudy (anonymous):

not sure but the answers are A. a[n]  = 80 + (n − 1) B. a[n]  = 5 (n − 1) C. a[n]  = 80 + 5 (n − 1) D. a[n]  = 80n E. None of the above

OpenStudy (amistre64):

we can start at a4, by subtracting 4 from each position a[4] = 95 a[7] = 110 -4 -4 ------------------- a0 = 95 a3 = 110

OpenStudy (amistre64):

since a arith seq is defined generically as: an = a0 + d(n-1) we can fill in the parts now a3 = 95 + d(3-1) = 110 ; to find d

OpenStudy (anonymous):

a+3d=95 a+6d=110 ----------- 3d=15 => d=5 => a=95-3*5=80 so a[n]=80+5(n-1)

OpenStudy (anonymous):

thank you both <3

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

good luck @laurenbae